1. 3(5x-9)-x=-12+4x

2. 20-6x=-4(2X-1)+2
please show the steps to the answer so that I can get an understanding of how to complete the problem. Thank you.

Answer by stanbon(75887)

1. 3(5x-9)-x=-12+4x
15x-27-x = -12+4x
14x - 27 =-12 + 4x
10x = 39
x = 3.9
---------------------
2. 20-6x=-4(2X-1)+2
20-6x = -80x+4+2
20-6x = -80x + 6
74x = -14
x = -7/37
-----------------
Cheers,
Stan H.